OpenStudy (anonymous):

would (3/4)*(1/x^(-1/2)) be (3/4)(x^(-1/2)) or (3/4)(x^(1/2))

hartnn (hartnn):

\(\dfrac{1}{x^{-n}}=x^n\)

OpenStudy (anonymous):

thats what i thought, but my answer sheet says that its x^-1/2

hartnn (hartnn):

whats the entire initial question ?

OpenStudy (anonymous):

convert \[\frac{ 3 }{ 4\sqrt{x} }\] to the form \[ax ^{n}\]

hartnn (hartnn):

\(\dfrac{1}{x^n}=x^{-n}\) so, \(\Large \dfrac{3}{4x^{1/2}}=\dfrac{3x^{-1/2}}{4}\)

hartnn (hartnn):

did u mis-intrepret \(\sqrt x \: as \: x^{-1/2}, \: its \: x^{1/2}\)

OpenStudy (anonymous):

no the answer steps show this: \[\frac{ 3 }{ 4\sqrt{x} }\] = \[\left( \frac{ 3 }{ 4 } \right)\left( \frac{ 1 }{ x ^{-\frac{ 1 }{ 2 }} } \right)\] = \[\frac{ 3 }{ 4 }x ^{-\frac{ 1 }{ 2 }}\]

hartnn (hartnn):

there is a typing mistake in 2nd step the denominator must have the exponent of x as +1/2 because \(\huge \sqrt x=x^{1/2}\)

OpenStudy (anonymous):

oh, so the third step is still correct isn't it?

hartnn (hartnn):

the final answer is still correct :)